# Amorphous Boron Nitride as an Ultrathin Copper Diffusion Barrier for Advanced Interconnects

### v3

@@ Abstract
This study focuses on amorphous boron nitride (-BN) as a novel diffusion barrier for advanced semiconductor technology, particularly addressing the critical challenge of copper diffusion in back-end-of-line (BEOL) interconnects. Owing to its ultralow dielectric constant and robust barrier properties, -BN is examined as an alternative to conventional low-k dielectrics. The investigation primarily employs theoretical modelling, using a Gaussian Approximation Potential, to simulate and understand the atomic-level interactions. This machine learning-based approach allows the performance of realistic simulations of amorphous structure of -BN, enabling the exploration of the impact of different film morphologies on barrier efficacy. Furthermore, we studied the electronic and optical properties of the films using a simple Tight-Binding model. In addition to the theoretical studies, we performed diffusion studies of copper through PECVD -BN on Si. The results from both the theoretical and experimental investigations highlight the potential of -BN as a highly effective diffusion barrier, suitable for integration in nanoelectronics. This research shows that -BN is a promising candidate for BEOL interconnects but also demonstrates the synergy of advanced computational models and experimental methods in material innovation for semiconductor applications.
@@ 1 Introduction
The continuous down-scaling of devices and architectures has been the major goal of the electronic industry for decades in order to increase performance, chip density, and reduce power consumption [1, 2, 3]. Interconnect technology is also experiencing significant challenges as devices are scaled down, with sustained efforts to decrease the resistance of metals and to lower the dielectric constant of intermetal dielectrics (IMDs), for optimised control of signal delays such as RC delay.
@@
The diffusion barrier that prevents the interaction between metals and IMD is also becoming thinner and is required to have excellent blocking properties. Ultimately, the technology aims for an almost “barrier-free” structure. Two-dimensional (2D) materials, such as and have recently garnered attention as potential barrier materials. Indeed, recent studies have shown that , when synthesized through Atomic Layer Deposition (ALD), can act as an effective Cu diffusion barrier, potentially reducing barrier thickness down to 1-2 nm[4, 5, 6]. Furthermore, the integration of dopants into such as Niobium (Nb) has been found to enhance its ability to block Cu atom diffusion, indicating its suitability for sub-5 nm technology nodes. However, while these 2D materials could act as efficient diffusion barriers, they also have a higher dielectric constant which would add to the RC delay.
@@
Hexagonal boron nitride (-BN), on the other hand, could be a better choice due to its low dielectric constant and higher bandgap. However, the practical application of these 2D materials is confronted with significant challenges. Integrating these materials into existing semiconductor manufacturing processes, especially developing a BEOL-compatible growth process for deposition on dielectrics remains a major challenge [7, 8, 9]. Moreover, the presence of defects in 2D materials, such as vacancy defects in hBN and , can lead to Cu accumulation, creating diffusion pathways that compromise the barrier’s effectiveness [10]. Amidst the ongoing quest for materials that can meet the stringent demands of advanced semiconductor technologies, amorphous boron nitride (-BN) emerges as a particularly compelling candidate. In contrast to h-BN, -BN does not suffer from dangling bonds and grain boundaries. In addition to its low temperature growth, -BN films can be uniformly synthesized on various substrates (such as metals and dielectrics) with a controllable thickness. This makes -BN a desirable dielectric material and diffusion barrier [11, 12]. Besides, -BN has been reported to have a low dielectric constant as a result of poor dipole alignment due to its amorphous nature [13]. While Hong et al.[2] reported an ultralow dielectric constant of 1.7 for 3 nm PECVD-grown -BN. Glavin et al.[12] pointed out that homonuclear bonds may open conductive pathways and increase the dielectric constant. They measured a device dielectric constant of atomic layer deposition (ALD)-grown -BN of 5.9 with a band gap of 4.5 eV [12]. Similarly, the dielectric constant of ALD grown -BN at 65 - 250 ºC is between 4.3 - 9.0 which increases with faster cooling. They observed that faster cooling leads to N deficiency and higher oxygen contamination, causing midgap states and higher dielectric constant [14]. Lin et al.[15] achieved a dielectric constant of 2.1 for sub-10 nm -BN films by optimising the growth process and using borazine as a precursor. Meanwhile, Abbas et al.[16] reported dielectric constants of 2.61 - 5.88 for sputtered -BN films with thicknesses ranging between 12 nm and 800 nm[16]. Due to their low dielectric constant, good mechanical properties, and thermal stability, -BN films have already been shown to act as efficient metal diffusion barriers. While Hong et al. [2] showed that a 3 nm -BN film can stop Co diffusion at 600 ºC, Kim et al. [17] found that 7-nm--BN film can stop Cu diffusion up to 500 ℃. Hence, -BN seems to be the ideal dielectric material due to its ultralow dielectric constant and great barrier properties. However, as shown in the literature [12, 15, 14, 18, 19], the properties and performance of -BN are determined by their morphology. Hence, ALD -BN with the correct composition could be an ideal dielectric material due to the potential low dielectric constant and excellent metal barrier properties.
@@
An exhaustive experimental investigation of the morphology and properties of -BN would be quite time consuming and expensive. On the other hand, atomistic simulations would allow us to study materials at the atomic level and can provide faster insights into the relation between morphology and key parameters of materials with moderate computational cost. Notwithstanding, implementing an accurate modelling of highly disordered materials such as -BN is quite a challenging task. Due to the lack of long-range order, the simulation box of amorphous materials should be quite large. This becomes quickly prohibitive for ab initio calculations. On the other hand, simulations using classical Molecular Dynamics (MD) with empirical potentials can deal with such large samples. However, due to simplifications and assumptions used to develop these interatomic potentials, classical MD cannot capture the diverse atomic environments of the amorphous material. These challenges can be overcome using machine learning-based approaches. Classical MD simulations with machine learning-based interatomic potential trained over ab initio calculations can deal with very large simulation boxes with ab initio accuracy [20, 21].
@@ 2.1 Molecular Dynamics Simulations
MD simulations use Newton’s law of motion to follow the trajectories of atoms through time. The accuracy and reliability of MD simulations largely depends on the accuracy of interatomic potentials. While there are several empirical potentials for -BN and Cu separately, there is no potential that takes both BN and Cu into account. Moreover, empirical potentials provide insufficient accuracy for such complex systems [20, 21, 22]. Machine learning based interatomic potentials that utilize a large dataset of quantum-mechanically computed reference data for structures, energies and forces are quite useful for disordered systems since they enable us to study large numbers of atoms with an accuracy comparable to DFT and a reasonable computing cost. Therefore, we trained a Gaussian Approximation Potential for BNCu systems that is based on DFT-generated energy and force data of a large database. The dataset contains isolated atoms, molecular compounds, crystalline structures, heterostructures and random structures. Details of training and validation of the GAP model are presented in A.
To generate -BN films with different qualities, we employed a melt-quench protocol. This protocol is an efficient way to generate amorphous structures and glasses in MD and has been used extensively by community [18, 19, 22]. In this protocol, an equal number of B and N atoms were placed in a cube simulation box which is periodic in all directions. However, we placed a 25 -thick vacuum region at both the top and bottom of the film to prevent boundary interactions and mimic thin films. After a minimization and an equilibration run at 5000 K for 20 ps with a timestep of 0.25 fs (same timestep used in all simulation), all films were cooled down with a constant cooling rate of 5 - 100 K/ps under NVT ensemble (constant volume-constant temperature) to systematically study the effect of cooling rates on film morphology, capturing variations in the degree of disorder characteristic of amorphous materials. Later all films are relaxed under NVE (constant volume-constant energy) and equilibrated under NVT ensemble for 20 ps each. Here, the initial density of each cube (before the addition of the vacuum regions) was fixed at approximately 2.1 , the experimentally reported value for -BN films [2]. The number of atoms in 3-nm, 5-nm and 7-nm thick films are respectively 3000, 13000 and 37000 to maintain the selected density. Upon generating the films, several morphological features are investigated to evaluate the quality of samples. We checked the presence of and atoms, homonuclear bonds, nanovoids, as well as the density of the samples. We generated 5 samples at each thickness with each cooling rate to get statistically meaningful data. A 3-nm thick -BN sample generated using this protocol is shown in Fig.1-a.
@@ 2.2 Tight Binding Model
in which is the system’s unit cell volume, is the reduced Planck constant, is the electron mass, is the polarization of the electric field, its angular frequency, the and are respectively the electronic eigenenergies and eigenstates of the TB Hamiltonian, with and running over the valence (occupied) and conduction (unoccupied) states, and is the momentum operator.
@@
It has been suggested [28] that -BN can exhibit midgap states due to structural disorder and bond characteristics. These midgap states may increase the dielectric constant drastically since they allow low energy transitions. However, these states appear mostly localized in real space due to the effects of disorder in the samples. To visualize the formation of midgap states and these localization effects, we also calculated the electronic density of states (DOS)333The DOS are computed using the same eigenenergies as the ones used for the dielectric function calculation, and with the same lorentzian broadening of . and inverse participation ratio (IPR) for all eigenstates of each film. The IPR provides a measurement of a state’s localization in real space. For an electronic eigenstate , we adopt the definition , where the sum runs over atomic positions and is the corresponding probability of presence of the electron at atomic site (obtained by summing the modulus squared of the projection of on each of its orbitals). Heuristically, if describes a state localized on a single atomic site, say , and . On the other hand, if describes a state delocalized over the whole system, the will be of order and therefore .
@@ 2.3 Experimental Studies
To experimentally demonstrate the barrier properties of -BN samples, we employed capacitively coupled plasma-chemical vapor deposition (CCP-CVD) to deposit -BN films. The Si substrate was placed on a substrate holder in the deposition chamber. The substrate was heated at a rate of 10 ºC/min and reached growth temperatures between 350 and 400 ºC. Following a 20-minute annealing process, a borazine precursor (, Gelest) was introduced into the chamber at a rate of 0.05 sccm with a hydrogen flow of 20 sccm. During the growth process, the plasma operated at a power of 20 W. The growth time was adjusted between 20 and 80 minutes depending on the thickness of the -BN film. After the growth was completed, the furnace was slowly cooled down to room temperature.
@@ 2.3.2 Experimental Cu Diffusion Barrier Test
We assessed the Cu diffusion barrier properties of the grown -BN films. For comparison with simulation results, 3-nm, 5-nm, and 7-nm thick -BN films on Si substrate were prepared and a 50-nm thick Cu film was deposited on -BN/Si by e-beam evaporation. Cu/-BN/Si samples were annealed at the temperature range of 400 ºC to 600 ºC for 30 minutes using a vacuum furnace. Subsequently, these samples were immersed in a Cu etchant (49-1, Trensene) for 20 minutes to remove the Cu films, followed by a deionized water rinse. to remove the excess etchant. Cu diffusion was evaluated through confirmation of the formation of copper silicide using X-ray diffraction (XRD) and scanning electron microscopy (SEM).
@@ 3.1 Film Generation and Morphological Features
After generating samples through the melt-quench protocol, we first focused on key morphological features to identify the structural differences between fast- and slow-cooled structures. It is known that the morphology of -BN, as well as other disordered samples, determines the properties and performance of a device. [12, 18, 15, 28]. Therefore, a thorough investigation of morphological features is vital. In order to the evaluate the “crystallinity vs. disorder” of the samples, we first computed the radial distribution function (RDF), , of -BN films as a function of cooling rate. Fig. 2 presents the average RDF of 3-nm thick films (since all films show the same trend, only the 3-nm thick films are presented). While all structures are characterized by a clear and recognizable first two peaks, no peak can be identified for distances larger than 3 , indicating the absence of long-range order in the samples regardless of the cooling protocol employed and showing the amorphous character of the films is not affected by the cooling rate. The first peaks of the slow-cooled samples are close to 1.44-1.45 , which is the bond distance of B-N bonds, as can be seen in the close-up image of the first peak of the RDF presented in the inset of Fig. 2. The broad peak in the RDF indicates a significant variation in bond lengths, reflecting a distribution between 1.35 and 1.55 , due to structural disorder. The peak becomes more broadened and shifted from 1.45 with higher cooling rates, indicating more disordered structures. Fig. 3 presents how several morphological features are changing due to the applied cooling rate. As shown in the figure, with higher cooling rates, the number of B-B bonds (which usually have a bond length of 1.6 - 1.8 ) increased rapidly. Also, we observed that -hybridized atoms, since they are mostly located around voids and pinholes, have quite strongly varying bond distances, which also contributes to the broadening observed in the RDF.
@@
Fig. 3 presents many interesting trends between cooling rates and morphological features. Fig.3-a shows that higher cooling rates increase the proportion of both B-B and N-N bonds for all thickness. However, the amount of increase in homonuclear bonds varies across film thicknesses. For any given cooling rate, thinner samples consistently have more homonuclear bonds than thicker ones. For instance, the portions of homonuclear bonds in 3-nm and 5-nm thick films generated at a cooling rate of 5 K/ps are approximately 10, while this rises to 30 at higher cooling rates. In contrast, for rapidly cooled (cooling rate 50 K/ps) the 7-nm thick sample, the proportion of homonuclear bonds remains below 20 . Fig. 3-b shows how the coordination number of -BN is influenced by applied cooling rates during film generation. Similarly to what we saw for homonuclear bonds, the effects of cooling rate on the local coordination numbers are more pronounced for thinner films. While all structures cooled at 5 K/ps have no -hybridized atoms (coordination number of 2), their proportion rapidly increases to 10 in 3-nm thick films at higher cooling rates. Similarly, for 5- and 7-nm thick films, the proportions of -hybridized atoms become and , respectively. Moreover, the number of -hybridized atoms (coordination number of 4) drops rapidly at all thicknesses as cooling rates increase. These changes in hybridization are also reflected in the average coordination number of the films, as shown in Fig. 3-c, where larger cooling rates result in a rapid decrease in average coordination for all thicknesses, most notably in the 3-nm thick films. While - and -hybridized fractions show opposing trends with the cooling rate, the average coordination number reflects the overall change in the local atomic environment. With larger cooling rates, the average coordination number of samples at all thicknesses drops quickly, with the largest change observed for 3 nm thick films. A parallel trend is also observed in density of the samples,444Like for the dielectric function, we used the difference in z-coordinate between the atoms of highest and lowest and the area of the periodic cell to estimate the film volume . as depicted in Fig. 3-d. The decrease in coordination number, as cooling rates increase, shows atoms getting far away from each other which leads lower density. While the density of films is for samples cooled with rate of 5 K/ps, the density drops even lower than 2 for 3-nm and 5-nm thick films at higher cooling rate. Meanwhile, even the 7-nm thick samples generated with the largest cooling rate still has an average density of . The significant reduction in density, combined with the increasing number of -hybridized atoms suggests that fast-cooled films have more porous structures. Fig. 11 presented in B shows some voids observed in fast-cooled (with 50 K/ps and 100 K/ps) 3-nm thick -BN films. Around these defects, we also observe a number of -hybridized atoms, which underlines the relationship between the existence of voids and -hybridized atoms.
@@
The observed reduction in homonuclear bonds and -hybridized atoms with increasing thickness, as shown in Fig. 3, suggests enhanced structural relaxation in thicker films. Interestingly, this improvement in structural order becomes more pronounced in the 7-nm films compared to the 3-nm and 5-nm films, where the differences are relatively smaller. This significant improvement in the 7-nm films reflects the ability to better accommodate strain and achieve a more ordered atomic arrangement as thickness increases. Similar behaviour has been reported in the literature for amorphous systems [35, 36, 37], where larger volumes facilitate relaxation processes and reduce disorder.

### v4

@@ Abstract
This study focuses on amorphous boron nitride (-BN) as a novel diffusion barrier for advanced semiconductor technology, particularly addressing the critical challenge of copper diffusion in back-end-of-line (BEOL) interconnects. Owing to its ultralow dielectric constant and robust barrier properties, -BN is examined as an alternative to conventional low-k dielectrics. The investigation primarily employs theoretical modelling, using a Gaussian Approximation Potential, to simulate and understand the atomic-level interactions. This machine learning-based approach allows the performance of realistic simulations of amorphous structure of -BN, enabling the exploration of the impact of different film morphologies on barrier efficacy. Furthermore, we studied the electronic and optical properties of the films using a simple Tight-Binding model. In addition to the theoretical studies, we performed diffusion studies of copper through PECVD -BN on Si. The results from both the theoretical and experimental investigations highlight the potential of -BN as a highly effective diffusion barrier, suitable for integration in nanoelectronics. This research shows that -BN is a promising candidate for BEOL interconnects combining the machine learned MD simulations with PECVD measurements.
@@ 1 Introduction
The continuous down-scaling of devices and architectures has been the major goal of the electronic industry for decades in order to increase performance, chip density, and reduce power consumption [1, 2, 3]. Interconnect technology is also experiencing significant challenges as devices are scaled down, with sustained efforts to decrease the resistance of metals and to lower the dielectric constant of intermetal dielectrics (IMDs), for optimised control of signal delays such as RC delay. The diffusion barrier that prevents the interaction between metals and IMD is also becoming thinner and is required to have excellent blocking properties. Ultimately, the technology aims to minimise barrier thickness in BEOL. Two-dimensional (2D) materials, such as and have recently garnered attention as potential barrier materials. Indeed, recent studies have shown that , when synthesized through Atomic Layer Deposition (ALD), can act as an effective Cu diffusion barrier, potentially reducing barrier thickness down to 1-2 nm[4, 5, 6]. Furthermore, the integration of dopants into such as Niobium (Nb) has been found to enhance its ability to block Cu atom diffusion, indicating its suitability for sub-5 nm technology nodes. However, while these 2D materials could act as efficient diffusion barriers, they also have a higher dielectric constant which would add to the RC delay.
@@
Hexagonal boron nitride (-BN), on the other hand, could be a better choice due to its low dielectric constant and higher bandgap. However, the practical application of these 2D materials is confronted with significant challenges. Integrating these materials into existing semiconductor manufacturing processes, especially developing a BEOL-compatible growth process for deposition on dielectrics remains a major challenge [7, 8, 9]. Moreover, the presence of defects in 2D materials, such as vacancy defects in hBN and , can lead to Cu accumulation, creating diffusion pathways that compromise the barrier’s effectiveness [10]. Amidst the ongoing quest for materials that can meet the stringent demands of advanced semiconductor technologies, amorphous boron nitride (-BN) emerges as a viable candidate. In contrast to h-BN, -BN does not suffer from dangling bonds and grain boundaries. In addition to its low temperature growth, -BN films can be uniformly synthesized on various substrates (such as metals and dielectrics) with a controllable thickness. This makes -BN a desirable dielectric material and diffusion barrier [11, 12]. Besides, -BN has been reported to have a low dielectric constant as a result of poor dipole alignment due to its amorphous nature [13]. While Hong et al.[2] reported an ultralow dielectric constant of 1.7 for 3 nm PECVD-grown -BN. Glavin et al.[12] pointed out that homonuclear bonds may open conductive pathways and increase the dielectric constant. They measured a device dielectric constant of atomic layer deposition (ALD)-grown -BN of 5.9 with a band gap of 4.5 eV [12]. Similarly, the dielectric constant of ALD grown -BN at 65 - 250 ºC is between 4.3 - 9.0 which increases with faster cooling. They observed that faster cooling leads to N deficiency and higher oxygen contamination, causing midgap states and higher dielectric constant [14]. Lin et al.[15] achieved a dielectric constant of 2.1 for sub-10 nm -BN films by optimising the growth process and using borazine as a precursor. Meanwhile, Abbas et al.[16] reported dielectric constants of 2.61 - 5.88 for sputtered -BN films with thicknesses ranging between 12 nm and 800 nm[16]. Due to their low dielectric constant, good mechanical properties, and thermal stability, -BN films have already been shown to act as efficient metal diffusion barriers. While Hong et al. [2] showed that a 3 nm -BN film can stop Co diffusion at 600 ºC, Kim et al. [17] found that 7-nm--BN film can stop Cu diffusion up to 500 ℃. Hence, -BN seems to be the ideal dielectric material due to its ultralow dielectric constant and great barrier properties. However, as shown in the literature [12, 15, 14, 18, 19], the properties and performance of -BN are determined by their morphology. Hence, ALD -BN with the correct composition could be an ideal dielectric material due to the potential low dielectric constant and excellent metal barrier properties.
@@
An exhaustive experimental investigation of the morphology and properties of -BN would be quite time consuming and expensive. On the other hand, atomistic simulations would allow us to study materials at the atomic level and can provide faster insights into the relation between morphology and key parameters of materials with moderate computational cost. Notwithstanding, implementing an accurate modelling of highly disordered materials such as -BN is quite a challenging task. Due to the lack of long-range order, the simulation box of amorphous materials should be quite large. This becomes quickly prohibitive for ab initio calculations. On the other hand, simulations using classical Molecular Dynamics (MD) with empirical potentials can deal with such large samples. However, due to simplifications and assumptions used to develop these interatomic potentials, classical MD cannot capture the diverse atomic environments of the amorphous material. These challenges can be overcome using machine learning-based approaches. Classical MD simulations with machine learning-based interatomic potential trained over ab initio calculations can deal with very large simulation boxes with ab initio accuracy [20, 21, 22, 23].
@@ 2.1 Molecular Dynamics Simulations
MD simulations use Newton’s law of motion to follow the trajectories of atoms through time. The accuracy and reliability of MD simulations largely depends on the accuracy of interatomic potentials. While there are several empirical potentials for -BN and Cu separately, there is no potential that takes both BN and Cu into account. Moreover, empirical potentials provide insufficient accuracy for such complex systems [20, 21, 24]. Machine learning based interatomic potentials that utilise a large dataset of quantum-mechanically computed reference data for structures, energies and forces are quite useful for disordered systems since they enable us to study large numbers of atoms with an accuracy comparable to DFT and a reasonable computing cost. Therefore, we trained a Gaussian Approximation Potential for BNCu systems that is based on DFT-generated energy and force data of a large database. The dataset contains isolated atoms, molecular compounds, crystalline structures, heterostructures and random structures. To ensure accuracy beyond nearest neighbours, the GAP employs SOAP descriptors with a cutoff rcut=7 , which spans first, second, and higher shells in -BN and include intermediate medium-range interactions in -BN [25]. Details of training and validation of the GAP model are presented in A.
To generate -BN films with different qualities, we employed a melt-quench protocol. This protocol is an efficient way to generate amorphous structures and glasses in MD and has been used extensively by community [18, 19, 24]. In this protocol, an equal number of B and N atoms were placed in a cube simulation box which is periodic in all directions. However, we placed a 25 -thick vacuum region at both the top and bottom of the film to prevent boundary interactions and mimic thin films. After a minimisation and an equilibration run at 5000 K for 20 ps with a timestep of 0.25 fs (same timestep used in all simulation), all films were cooled down with a constant cooling rate (CR) of 5 - 100 K/ps under NVT ensemble (constant volume-constant temperature) to systematically study the effect of cooling rates on film morphology, capturing variations in the degree of disorder characteristic of amorphous materials. Later all films are relaxed under NVE (constant volume-constant energy) and equilibrated under NVT ensemble for 20 ps each. Here, the initial density of each cube (before the addition of the vacuum regions) was fixed at approximately 2.1 , the experimentally reported value for -BN films [2]. The number of atoms in 3-nm, 5-nm and 7-nm thick films are respectively 3000, 13000 and 37000 to maintain the selected density. Upon generating the films, several morphological features are investigated to evaluate the quality of samples. We checked the presence of and atoms, homonuclear bonds, nanovoids, as well as the density of the samples. We generated 5 samples at each thickness with each cooling rate to get statistically meaningful data. A 3-nm thick -BN sample generated using this protocol is shown in Fig.1-a.
@@ 2.2 Tight Binding Model
in which is the system’s unit cell volume, is the reduced Planck constant, is the electron mass, is the polarisation of the electric field, its angular frequency, the and are respectively the electronic eigenenergies and eigenstates of the TB Hamiltonian, with and running over the valence (occupied) and conduction (unoccupied) states, and is the momentum operator. For clarity, we note that such a frozen atoms calculation excludes the vibrational contributions to the dielectric function.
@@
It has been suggested [31] that -BN can exhibit midgap states due to structural disorder and bond characteristics. These midgap states may increase the dielectric constant drastically since they allow low energy transitions. However, these states appear mostly localised in real space due the disordered nature of the samples in the samples. To visualise the formation of midgap states and their localisation effects, we calculated the electronic density of states (DOS)333The DOS are computed using the same eigenenergies as the ones used for the dielectric function calculation, and with the same lorentzian broadening of . and the inverse participation ratio (IPR) for all eigenstates of each film. The IPR provides a measurement of a state’s localisation in real space. For an electronic eigenstate , we adopt the definition , where the sum runs over atomic positions and is the corresponding probability of presence of the electron at atomic site (obtained by summing the modulus squared of the projection of on each of its orbitals). Heuristically, if describes a state localised on a single atomic site, say , and . On the other hand, if describes a state delocalised over the whole system, the will be of order and therefore .
@@ 2.3 Experimental Studies
To experimentally demonstrate the barrier properties of -BN samples, we employed capacitively coupled plasma-chemical vapour deposition (CCP-CVD) to deposit -BN films. The Si substrate was placed on a substrate holder in the deposition chamber. The substrate was heated at a rate of 10 ºC/min and reached growth temperatures between 350 and 400 ºC. Following a 20-minute annealing process, a borazine precursor (, Gelest) was introduced into the chamber at a rate of 0.05 sccm with a hydrogen flow of 20 sccm. During the growth process, the plasma operated at a power of 20 W. The growth time was adjusted between 20 and 80 minutes depending on the thickness of the -BN film. After the growth was completed, the furnace was slowly cooled down to room temperature.
@@ 2.3.2 Experimental Cu Diffusion Barrier Test
We assessed the Cu diffusion barrier properties of the grown -BN films. For comparison with simulation results, 3-nm, 5-nm, and 7-nm thick -BN films on Si substrate were prepared and a 50-nm thick Cu film was deposited on -BN/Si by e-beam evapouration. Cu/-BN/Si samples were annealed at the temperature range of 400 ºC to 600 ºC for 30 minutes using a vacuum furnace. Subsequently, these samples were immersed in a Cu etchant (49-1, Trensene) for 20 minutes to remove the Cu films, followed by a deionized water rinse. to remove the excess etchant. Cu diffusion was evaluated through confirmation of the formation of copper silicide using X-ray diffraction (XRD) and scanning electron microscopy (SEM).
@@ 3.1 Film Generation and Morphological Features
After generating samples through the melt-quench protocol, we focused on key morphological features to identify the structural differences between fast- and slow-cooled structures. The morphology of -BN, as with other disordered materials, critically influences material properties and device performance [12, 18, 15, 31]. A thorough investigation of morphological features is therefore vital. To evaluate the “crystallinity vs. disorder” of the samples, we first computed the radial distribution function (RDF), , of -BN films as a function of cooling rate. Fig. 2 presents the average RDF of 3-nm thick films. As all thicknesses show the same trend, we only present the 3-nm thick films here, RDF of 5- and 7- nm thick films is presented in Fig. 11 in B. While all structures are characterised by a clear and recognisable first two peaks, no peak is identified for distances larger than 3 , indicating the absence of long-range order in the samples regardless of the cooling protocol employed and showing the amorphous character of the films is not affected by the cooling rate. The first peaks of the slow-cooled samples are close to 1.44-1.45 , which is the bond distance of B-N bonds, as can be seen in the close-up image of the first peak of the RDF presented in the inset of Fig. 2. The broad peak in the RDF indicates a significant variation in bond lengths, reflecting a distribution between 1.35 - 1.55 , and is consistent with structural disorder.The peak broadens further and shifts away from 1.45 with higher cooling rates, indicating more disordered structures. Similar behaviour is observed in the 5-nm and 7-nm films as shown in Figs. 11 in B. With increasing film thickness, the first-neighbour peak in the RDF sharpens slightly, indicating improved short-range order. However, no additional peaks emerge beyond , confirming the lack of long-range order even in the thickest films. This is consistent with the absence of medium-range features in the corresponding structure factors (Fig. 12 in B). shows no features associated with medium-range order, and no first sharp diffraction peak is present [25]. While the intensity of the main peak in increases modestly with thickness, no higher-order peaks appear, indicating that the structural network remains amorphous regardless of thickness. Fig. 3 presents how several morphological features change with the applied cooling rate. As shown in the figure, with higher cooling rates, the number of B-B bonds (which usually have a bond length of 1.6 - 1.8 ) increased rapidly. In addition, we observed that -hybridised atoms, since they are mostly located around voids and pinholes, have quite strongly varying bond distances, which also contributes to the broadening observed in the RDF.
@@
Fig. 3 summarises trends between cooling rate and morphological features. Fig. 3-a shows that higher cooling rates increase the proportion of both B-B and N-N bonds for all thicknesses. However, the increase in homonuclear bonds varies with film thickness. At any given cooling rate, thinner samples consistently contain more homonuclear bonds than thicker ones. For instance, the fraction of homonuclear bonds in 3-nm and 5-nm thick films generated at a cooling rate of 5 K/ps are approximately 10, rising to 30 at higher cooling rates, whereas for rapidly cooled (cooling rate 50 K/ps) the 7-nm thick sample, the proportion of homonuclear bonds remains below 20 . Fig. 3-b shows how the coordination number of -BN is influenced by applied cooling rate during film generation. As with the homonuclear fraction, the effect of cooling rate on local coordination is more evident in thinner films. While all structures cooled at 5 K/ps have no -hybridised atoms (coordination number of 2), their proportion rapidly increases to 10 in 3-nm thick films at higher cooling rates. Similarly, for 5- and 7-nm thick films, the fractions of -hybridised atoms become and , respectively. Moreover, the number of -hybridised atoms (coordination number of 4) drops rapidly at all thicknesses as cooling rate increases. These changes in hybridisation are reflected in the average coordination number (Fig. 3-c), larger cooling rates lead to a rapid decrease in average coordination at all thicknesses, most notably in the 3-nm films. While - and -hybridised fractions show opposing trends with the cooling rate, the average coordination number reflects the overall change in the local atomic environment. With larger cooling rates, the average coordination number of samples at all thicknesses drops quickly, with the largest change observed for 3 nm thick films. A parallel trend is also observed in mass density of the samples. As for the dielectric function, we estimated the film volume as using the difference in -coordinate between the atoms of highest and lowest and the area of the periodic cell, as depicted in Fig. 3-d. as depicted in Fig. 3-d. The decrease in coordination number, as cooling rates increase, indicates increased interatomic separation, which lowers the density. While the density of films is for samples cooled at 5 K/ps, the density drops below 2 for 3-nm and 5-nm thick films at higher cooling rate. Meanwhile, even the 7-nm thick samples generated with the largest cooling rate still have an average density of . The significant reduction in density, together with the increase in -hybridised atoms suggests that fast-cooled films have more porous structures.Fig. 13, presented in B, shows voids in fast-cooled (50 K/ps and 100 K/ps) 3-nm -BN films. Around these defects we also observe -hybridised atoms, emphasise the relationship between voids and environments.
@@
With increasing thickness, the reduction in homonuclear bonds and -hybridised atoms shown in Fig. 3 indicates further structural relaxation in thicker films. This improvement in structural order is most pronounced in the 7-nm films, whereas the 3-nm and 5-nm films show smaller changes. The behaviour in the 7-nm films suggests a greater capacity to accommodate strain and to attain a more ordered atomic arrangement as thickness increases. Similar behaviour has been reported for amorphous systems [38, 39, 40], where larger volumes facilitate relaxation processes and reduce disorder. Although thickness favours relaxation, larger thicknesses during growth can facilitate ring alignment and the emergence of turbostratic or nanocrystalline h-BN [17]. This competing crystallisation pathway can modify dielectric response and diffusion-barrier performance.